OpenStudy (anonymous):

PLEASE HELP!!! Two students, Ann and Max, factored the trinomial 4x2 − 6x − 4. Ann factored it as 2(x − 2)(2x + 1) and Max factored it as (x − 2)(4x + 2). Indicate which student factored the trinomial completely and which student did not, and explain why.

OpenStudy (anonymous):

Ann factored it completely, as she factored out a term that Max left in his factorization.

OpenStudy (anonymous):

Ann factored it completely, as she factored out a term that Max left in his factorization.

OpenStudy (anonymous):

can you tell me how you got that?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@pooja195 @Luigi0210 @phi @whpalmer4

OpenStudy (tkhunny):

You should "Factor It Completely" and see who agrees with you.

OpenStudy (anonymous):

how do i factor it @tkhunny ?

OpenStudy (whpalmer4):

Max factored it as \[(x − 2)(4x + 2)\] Look at the two terms there: can you factor out anything from \[(x-2)\]? How about from \[(4x+2)\]?

OpenStudy (whpalmer4):

Put another way, can you write down an equation for a multiplication of 2 terms that will give you one of those quantities in parentheses? If you can, than that quantity can be factored again. As an example, say we have \[(3x+6)\]Is it possible to write a multiplication where the answer is \(3x+6\)? How about\[3(x+2) = 3*x + 3*2 = 3x + 6\]That means if we see \(3x+6\), we can factor it as \(3(x+2)\) because \(3(x+2) = 3x+6\)